OpenStudy (anonymous):

The equation cos(pi/6-x) is equal to _____?

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \cos(π/6~~-~~x) = }\) \(\large\color{midnightblue}{ \cos(30~~-~~x) = }\) \(\large\color{midnightblue}{ \cos(30) \cos(x)+\sin(30) \sin(x) = }\) \(\large\color{midnightblue}{ (\sqrt{3}~/~2)~ \cos(x)+(1~/~2)~ \sin(x) }\) \(\large\color{midnightblue}{ \sqrt{3}(1~/~2)~ \cos(x)+(1~/~2)~ \sin(x) }\) \(\large\color{midnightblue}{ (1~/~2)~[~ \sqrt{3}~ \cos(x)+~ \sin(x)~] }\)